OpenStudy (anonymous):

integrate from 1to 2---- (x^3-2x^2+3)/x^2 please help me simplify and choose u

OpenStudy (kropot72):

To simplify, just divide each term in the numerator by x^2.

OpenStudy (jamesj):

Do you know how to integrate x^n ?

OpenStudy (jamesj):

\[ (x^3-2x^2+3)/x^2 = x^3/x^2 - 2 x^2/x^2 + 3/x^2 = x - 2 + 3/x^2\] Now integrate each of these terms on the right.